# Inference-Time Text-to-Video Alignment with Diffusion Latent Beam Search

### v2

@@ Abstract
The remarkable progress in text-to-video diffusion models enables photorealistic generations, although the contents of the generated video often include unnatural movement or deformation, reverse playback, and motionless scenes. Recently, an alignment problem has attracted huge attention, where we steer the output of diffusion models based on some quantity on the goodness of the content. Because there is a large room for improvement of perceptual quality along the frame direction, we should address which metrics we should optimize and how we can optimize them in the video generation. In this paper, we propose diffusion latent beam search with lookahead estimator, which can select a better diffusion latent to maximize a given alignment reward, at inference time. We then point out that the improvement of perceptual video quality considering the alignment to prompts requires reward calibration by weighting existing metrics. This is because when humans or vision language models evaluate outputs, many previous metrics to quantify the naturalness of video do not always correlate with evaluation. We demonstrate that our method improves the perceptual quality evaluated on the calibrated reward, VLMs, and human assessment, without model parameter update, and outputs the best generation compared to greedy search and best-of-N sampling under much more efficient computational cost. The experiments highlight that our method is beneficial to many capable generative models, and provide a practical guideline that we should prioritize the inference-time compute allocation into lookahead steps for reward estimation over search budget or denoising steps.  [^1] Website: https://sites.google.com/view/t2v-dlbs  [^2] Code: https://github.com/shim0114/T2V-Diffusion-Search
@@ 1 Introduction
In this paper, we propose Diffusion Latent Beam Search (DLBS) with lookahead estimator, an inference-time search over the reverse process (Figure 1; Left), which can select a better diffusion latent to maximize a given alignment reward. A lookahead estimator reduces the noise in the reward estimate, and a beam search robustly explores the latent paths, avoiding inaccurate $\operatorname{argmax}$ operation.
We then point out that the improvement of perceptual video quality considering the alignment to prompts requires reward calibration of existing metrics [12]. When we evaluate outputs by capable vision language models [13, 14] or human raters, many previous metrics to quantify the naturalness of video do not always correlate with each other, and optimal reward design to measure perceptual quality highly depends on the degree of dynamic descriptions in evaluation prompts. We design a weighted linear combination of multiple metrics, which is calibrated to perceptual quality and improves the correlation with preference.
We demonstrate that DLBS can induce high-quality outputs based on the calibrated reward, AI and human feedback (Figure 2), without model parameter update, and realize the best generation under much more efficient computational cost compared to greedy search [15, 11] and best-of-N sampling [16, 17]. The experiments also highlight that our method is beneficial to many SoTA models (e.g., Latte [18], CogVideoX [19], and Wan 2.1 [20]), and provide a practical guideline on that we should prioritize the compute allocation into lookahead steps for reward estimate more than search budget or denoising steps, at an inference time.
@@ 2 Preliminaries
Latent Diffusion Models   Latent diffusion models [21, 18] are a special class of diffusion probabilistic models [22, 23], and popular choices for high-resolution text-to-video generation [24, 25, 4], which considers the diffusion process in embedding space. Let $\textbf{x}_{0}$ be a video and encode it as $\textbf{z}_{0}=\text{Enc}(\textbf{x}_{0})$ using VAE [26]. Continuous-time forward diffusion process can be modeled as a solution to a stochastic differential equation (SDE) [27]: $d\textbf{z}=\textbf{f}(\textbf{z},t)dt+g(t)d\textbf{w},$ where $_{0}\sim p_{0}(\textbf{z})$ z ) is the latent as initial condition while $p_{t}(\textbf{z})$ is the marginal distribution of $\textbf{z}_{t}$, $\textbf{f}:\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R}^{d}$ is the drift coefficient, $g:\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient, and $\textbf{w}\in\mathbb{R}^{d}$ is $d$-dimensional standard Wiener process. $\textbf{f}(\cdot,\cdot)$ and $g(\cdot)$ are designed appropriately for the marginal distribution to reach $p_{T}(\textbf{z})\approx\mathcal{N}(0,\textbf{I})$ as $t\rightarrow T$ [28]. Reverse diffusion process generates samples $\textbf{z}_{0}$ through the following reverse-time SDE: $d\textbf{z}=[\textbf{f}(\textbf{z},t)-g(t)^{2}\nabla_{\textbf{z}}\log p_{t}(\textbf{z})]dt+g(t)d\bar{\textbf{w}},$ where $dt$ here is an infinitesimal negative time step from $T$ to 00 and $\bar{\textbf{w}}\in\mathbb{R}^{d}$ is a standard reverse-time Wiener process. We start this with $\textbf{z}_{T}\sim\mathcal{N}(0,\textbf{I})$. This SDE induces the marginal distribution on the data $p^{\text{pre}}(\textbf{z})$ (i.e., pre-trained diffusion models). While we omit the notation for simplicity, we consider the text-to-video generation problem, where the diffusion process is conditioned on text prompts c.
@@
[…]
@@
[…]
@@
Figure 2: Qualitative evaluation among different search methods. We test prompts from DEVIL-high (above) and DEVIL-static (below). DLBS-LA generates more dynamic and prompt-aligned videos than GS or BoN.
@@ 3.1 A Unified View on Practical Approximations
1:  $\textbf{z}^{1}_{T},\cdots,\textbf{z}^{B}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$   $\triangleright$ Initial $B$ beams
2:  for $t=T$ to $1$ do
@@
15:         $\triangleright$ Search $B$ higher-reward beams from $KB$ latents
@@
Note that the current diffusion latents $\boldsymbol{\nu}_{t}$ and posterior mean $\hat{\boldsymbol{\nu}}_{T|t}$ are sampled by following the standard Wiener process $d\textbf{w}$. This approximation is leveraged for inference-time alignment via greedy search [11, 39] or SMC [41] of diffusion latents. However, greedy search can result in sub-optimal generation affected by inaccurate reward estimate $r^{\prime}(\hat{\boldsymbol{\nu}}_{T|t})$ due to its noisy input. Moreover, it can be challenging to obtain an accurate density ratio term required in SMC for a high-dimensional domain such as video generation.
@@ 3.2 Mitigating Approximation Errors via Beam Search
Existing practical algorithms based on these three approximations, such as greedy search [11, 39], fall into sub-optimal generation due to the erroneous reward evaluation with a noisy estimate of the posterior mean [36], and $\operatorname{argmax}$ operator based on them. To resolve the error accumulation, we propose a simple yet robust modification, diffusion latent beam search (DLBS) with lookahead estimator. To clearly describe the practical implementation, we use the notation of a discrete-time diffusion process in the rest of the section (see Appendix D for the continuous-time diffusion process).
@@
2:  $\tilde{t}(s)\in\{t-1,\dots,\lfloor\frac{s}{T^{\prime}}(t-1)\rfloor,\dots,\lfloor\frac{1}{T^{\prime}}(t-1)\rfloor,0\}$
@@
4:  $\textbf{z}_{\tilde{t}(T^{\prime})}:=\textbf{z}_{t-1}$
5:  $\tilde{\textbf{z}}_{0|\tilde{t}(T^{\prime})}=\frac{\textbf{z}_{\tilde{t}(T^{\prime})}-\sqrt{1-\tilde{\alpha}_{T^{\prime}}}\epsilon_{\theta}(\textbf{z}_{\tilde{t}(T^{\prime})})}{\sqrt{\tilde{\alpha}_{T^{\prime}}}}$
6:  for $s=T^{\prime}$ to $1$ do
@@
10:  return: $(\textbf{z}_{t-1},\tilde{\textbf{z}}_{0|\tilde{t}(0)})$triangleright Latent and LA estimator
Lookahead Estimator   The other source of approximation errors than the $\operatorname{argmax}$ operator is a one-step estimator of the posterior mean $\hat{\mathbf{z}}_{0|t-1}$ from Tweedie’s formula, which is still noisy, especially in earlier time steps, and leads to inaccurate reward evaluation. To reduce errors in reward evaluation, we propose a lookahead (LA) estimator $\tilde{\mathbf{z}}_{0|\tilde{t}(0)}$, which is estimated by running $T^{\prime}$-step deterministic DDIM $(1<T^{\prime}<<T)$ while equally interpolating the rest of time steps from the current latent $\mathbf{z}_{t-1}$ to $\mathbf{z}_{0}$ (Algorithm 2). While requiring additional denoising steps, its cost is almost the same as naive DLBS because most computational costs come from when we decode $\textbf{z}_{0}$ (i.e., reward evaluation). Theoretically, enlarging the lookahead steps $T^{\prime}$ monotonically tightens the upper bound on the reward-approximation error (see Appendix E). Empirically, a modest horizon ($T^{\prime}=2,3,6$) delivers substantial search improvements (Figure 8; Left), but the marginal gains saturate, so pushing $T^{\prime}$ further yields little additional benefit (see Appendix L.4).
@@ 4 Calibrating Reward to Preference Feedback
Human evaluation is one of the most valuable assessments for generative models, yet gathering human feedback at scale is prohibitively costly. A practical approach to reduce the time and cost is to leverage AI feedback from VLMs [43], which has been shown to modestly align with human judgment on video quality [44, 45] (see Appendix J). In this work, we assume that the VLM evaluation works as an oracle, and we align model outputs with the preference of VLMs, which is reasonable due to their capability and the cost to be saved. Our qualitative and quantitative evaluations also confirm that the high-rated video by VLMs is generally good for us.
However, because alignment via inference-time search requires massive reward evaluation queries, we still need to build more tractable proxy rewards that do not rely on humans or external VLM APIs. The question here is what metrics for perceptual video quality can improve the feedback from VLMs. Because the criteria of videos preferred by humans are multi-objective, maximizing a single metric may lead to undesirable generation due to over-optimization. For instance, focusing exclusively on temporal consistency or frame-by-frame quality metrics can unintentionally reduce the video’s motion magnitude. In this section, we first review the possible video quality metrics (Section 4.1), evaluate the Pearson correlation between these and VLM feedback score, and then propose a reward calibration (Section 4.2), aiming to align the existing video rewards to VLMs by considering their weighted linear combination through the brute-force search of coefficients.
Figure 3: 2D-histogram and correlation between reward functions for perceptual video quality [12] and AI feedback from Gemini [14]. A single reward (e.g., subject consistency; blue) is often not aligned well with a preference from Gemini, which happens for all the prompt sets with different degrees of dynamics. The calibrated reward, a linear combination of perceptual metrics via brute-force search (green), achieves the best Pearson correlation coefficient in all settings (statistically significant with $p<0.01$).
@@ 4.1 Metric Reward for Perceptual Video Quality
Figure 4: The coefficient of calibrated reward $w_{i}$ with feedback from Gemini. Each set of prompts depicting different degree of dynamics requires different mixture of perceptual video qualities.
@@
Experimental Setup  We leverage Gemini-1.5 [14] and GPT-4o [13] as automated raters for generated videos. We provide a prompt and generated video as inputs, instructing VLMs to assign discrete scores (from 1 to 10) based on overall visual quality (e.g., clarity, resolution, brightness, and aesthetic appeal), the appropriateness of motion for either static or dynamic scenes, the smoothness and consistency of shapes and motions, and the degree of alignment with a prompt (see Appendix H).
We select four prompt sets from two distinct datasets (see Appendix F). DEVIL [52] classifies its prompts into five categories depending on the degree of movement, each further divided by subject type (e.g., cat, horse, truck, nature, etc.). We focus on three of five categories (high, medium, and static) and select one prompt randomly from each subject-subdivision within a chosen category. We also draw 30 random captions from the test split of MSRVTT [53], widely used as a video benchmark.
@@ 4.2 Correlation and Reward Calibration
Figure 3 illustrates the 2D-histogram and corresponding correlation between each metric and feedback from Gemini. Relying on a single metric often yields low correlation, which supports the multifaceted nature of perceptual video quality. See Appendix I.1 for further results, where we can see that the relative importance of each metric depends on the movement in the prompts; in highly dynamic DEVIL-high, the dynamic degree correlates more strongly with VLMs than consistency metrics. Conversely, subject consistency and motion smoothness play more prominent roles in less-dynamic DEVIL-medium or DEVIL-static. Because the aesthetic score focuses on frame-by-frame visual quality, it tends to correlate strongly with VLM in low-motion scenarios. Rapid movements and frequent transitions often introduce motion blur or abrupt changes in composition, reducing the frame-level aesthetic quality and thus weakening its correlation with VLMs.
Reward calibration, a weighted linear combination of these metrics, yields the highest correlation with Gemini (Figure 3, green). We select the best coefficients among brute-force candidates, based on the correlation with Gemini, for each set of prompts that have a different description of the degree of movement (Figure 4). The dynamic degree and motion smoothness exhibit pronounced variation among DEVIL-high, -medium, and -static, and the weights of other factors, such as text-video consistency, subject consistency, and aesthetic considerations, are also adjusted accordingly. In addition, Appendix K presents results from best-of-64 sampling with a single metric or calibrated reward, where a single metric often leads to over-optimization. This highlights the importance of reward calibration, appropriately weighting multiple criteria, as aligning with a request in the prompt.

### v3

@@ Abstract
The remarkable progress in text-to-video diffusion models enables the generation of photorealistic videos, although the content of these generated videos often includes unnatural movement or deformation, reverse playback, and motionless scenes. Recently, an alignment problem has attracted huge attention, where we steer the output of diffusion models based on some measure of the content’s goodness. Because there is a large room for improvement of perceptual quality along the frame direction, we should address which metrics we should optimize and how we can optimize them in the video generation. In this paper, we propose diffusion latent beam search with lookahead estimator, which can select a better diffusion latent to maximize a given alignment reward at inference time. We then point out that improving perceptual video quality with respect to alignment to prompts requires reward calibration by weighting existing metrics. This is because when humans or vision language models evaluate outputs, many previous metrics to quantify the naturalness of video do not always correlate with the evaluation. We demonstrate that our method improves the perceptual quality evaluated on the calibrated reward, VLMs, and human assessment, without model parameter update, and outputs the best generation compared to greedy search and best-of-N sampling under much more efficient computational cost. The experiments highlight that our method is beneficial to many capable generative models, and provide a practical guideline: we should prioritize the inference-time compute allocation into enabling the lookahead estimator and increasing the search budget, rather than expanding the denoising steps.  [^1] Website: https://sites.google.com/view/t2v-dlbs [^2] Code: https://github.com/shim0114/T2V-Diffusion-Search
@@ 1 Introduction
In this paper, we propose Diffusion Latent Beam Search (DLBS) with lookahead estimator, an inference-time search over the reverse process (Figure 1; Left), which can select a better diffusion latent to maximize a given alignment reward. A lookahead estimator reduces the noise in the reward estimate, and a beam search robustly explores the latent paths, avoiding inaccurate $\operatorname{argmax}$ operations.
We then point out that the improvement of perceptual video quality, considering the alignment to prompts, requires reward calibration of existing metrics [12]. When evaluating outputs using capable vision language models [13, 14] or human raters, many previous metrics for quantifying video naturalness do not always correlate with them. Optimal reward design for measuring perceptual quality highly depends on the degree of dynamics described in evaluation prompts. We design a weighted linear combination of multiple metrics, which is calibrated to perceptual quality and improves the correlation with VLM/human preference.
We demonstrate that DLBS can induce high-quality outputs based on the calibrated reward, AI, and human feedback (Figure 2), without model parameter update, and realize the best generation under much more efficient computational cost compared to greedy search [15, 11] and best-of-N sampling [16, 17]. The experiments also highlight that our method is beneficial to many SoTA models (e.g., Latte [18], CogVideoX [19], and Wan 2.1 [20]), and provide a practical guideline that we should prioritize the inference-time compute allocation into enabling the lookahead estimator and increasing the search budget, rather than expanding the denoising steps.
@@ 2 Preliminaries
Latent Diffusion Models   Latent diffusion models [21, 18] are a special class of diffusion probabilistic models [22, 23], and popular choices for high-resolution text-to-video generation [24, 25, 4], which considers the diffusion process in embedding space. Let $\textbf{x}_{0}$ be a video and encode it as $\textbf{z}_{0}=\text{Enc}(\textbf{x}_{0})$ using VAE [26]. Continuous-time forward diffusion process can be modeled as a solution to a stochastic differential equation (SDE) [27]: $d\textbf{z}=\textbf{f}(\textbf{z},t)dt+g(t)d\textbf{w},$ where $\textbf{z}_{0}\sim p_{0}(\textbf{z})$ is the latent as initial condition while $p_{t}(\textbf{z})$ is the marginal distribution of $\textbf{z}_{t}$, $\textbf{f}:\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R}^{d}$ is the drift coefficient, $g:\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient, and $\textbf{w}\in\mathbb{R}^{d}$ is $d$-dimensional standard Wiener process. $\textbf{f}(\cdot,\cdot)$ and $g(\cdot)$ are designed appropriately for the marginal distribution to reach $p_{T}(\textbf{z})\approx\mathcal{N}(0,\textbf{I})$ as $t\rightarrow T$ [28]. Reverse diffusion process generates samples $\textbf{z}_{0}$ through the following reverse-time SDE: $d\textbf{z}=[\textbf{f}(\textbf{z},t)-g(t)^{2}\nabla_{\textbf{z}}\log p_{t}(\textbf{z})]dt+g(t)d\bar{\textbf{w}},$ where $dt$ here is an infinitesimal negative time step from $T$ to 0 and $\bar{\textbf{w}}\in\mathbb{R}^{d}$ is a standard reverse-time Wiener process. We start this with $\textbf{z}_{T}\sim\mathcal{N}(0,\textbf{I})$. This SDE induces the marginal distribution on the data $p^{\text{pre}}(\textbf{z})$ (i.e., pre-trained diffusion models). While we omit the notation for simplicity, we consider the text-to-video generation problem, where the diffusion process is conditioned on text prompts c.
@@
Figure 2: Comparison of text-to-video results between DLBS-LA, base models, and other sampling methods on SoTA models (Latte [18], CogVideoX [19], and Wan 2.1 [20]). DLBS-LA produces more dynamic, natural, and prompt-aligned videos than all baselines.
@@
$\textbf{u}^{*}=\underset{\textbf{u}}{\operatorname{argmax}}~\mathbb{E}\left[r^{\prime}(\boldsymbol{\nu}_{T})-\frac{\lambda}{2}\int_{t=0}^{T}\frac{\|\textbf{u}(\boldsymbol{\nu}_{t},t)\|^{2}}{g(t)^{2}}dt\right]$ (1)
@@
$\begin{split}v^{*}_{t}(\boldsymbol{\nu})&=\mathbb{E}_{p^{*}}\left[r^{\prime}(\boldsymbol{\nu}_{T})-\frac{\lambda}{2}\int_{s=t}^{T}\frac{\|\textbf{u}(\boldsymbol{\nu}_{s},s)\|^{2}}{g(s)^{2}}ds|\boldsymbol{\nu}_{t}=\boldsymbol{\nu}\right],\end{split}$ (2)
@@ 3.1 A Unified View on Practical Approximations
[…]
2:for $t=T$ to 11 do
@@
15:    $\triangleright$ Search $B$ higher-reward beams from KBKB latents
@@
Note that the current diffusion latents $\boldsymbol{\nu}_{t}$ and posterior mean $\hat{\boldsymbol{\nu}}_{T|t}$ are sampled by following the standard Wiener process $d\textbf{w}$. This approximation is leveraged for inference-time alignment via greedy search [11, 39] or SMC [41] of diffusion latents. However, greedy search can result in sub-optimal generation affected by inaccurate reward estimate $r^{\prime}(\hat{\boldsymbol{\nu}}_{T|t})$ due to its noisy input. Moreover, it can be challenging to obtain an accurate density ratio term required in SMC for a high-dimensional domain, such as video generation.
@@ 3.2 Mitigating Approximation Errors via Beam Search
Existing practical algorithms based on these three approximations, such as greedy search [11, 39], fall into sub-optimal generation due to the erroneous reward evaluation with a noisy estimate of the posterior mean [36], and $\operatorname{argmax}$ operator based on them. To resolve the error accumulation, we propose a simple yet robust modification, diffusion latent beam search (DLBS) with lookahead estimator. To clearly describe the practical implementation, we use the notation of a discrete-time diffusion process in the rest of the section (see Appendix E for the continuous-time diffusion process).
@@
[…]
@@
[…]
6:for $s=T^{\prime}$ to 11 do
@@
10:return: $(\textbf{z}_{t-1},\tilde{\textbf{z}}_{0|\tilde{t}(0)})$ $\triangleright$ Latent and LA estimator
Lookahead Estimator   The other source of approximation errors than the $\operatorname{argmax}$ operator is a one-step estimator of the posterior mean $\hat{\mathbf{z}}_{0|t-1}$ from Tweedie’s formula, which is still noisy, especially in earlier time steps, and leads to inaccurate reward evaluation. To reduce errors in reward evaluation, we propose a lookahead (LA) estimator $\tilde{\mathbf{z}}_{0|\tilde{t}(0)}$, which is estimated by running $T^{\prime}$-step deterministic DDIM $(1<T^{\prime}<<T)$ while equally interpolating the rest of time steps from the current latent $\mathbf{z}_{t-1}$ to $\mathbf{z}_{0}$ (Algorithm 2). While requiring additional denoising steps, its cost is almost the same as naive DLBS because most computational costs come from when we decode $\textbf{z}_{0}$ (i.e., reward evaluation). Theoretically, enlarging the lookahead steps $T^{\prime}$ monotonically tightens the upper bound on the reward-approximation error (see Appendix F). Empirically, a modest horizon ($T^{\prime}=2,3,6$) delivers substantial search improvements (Figure 9; Left), but the marginal gains saturate, so pushing $T^{\prime}$ further yields little additional benefit (see Appendix M.4).
@@ 4 Calibrating Reward to Preference Feedback
Human evaluation is one of the most valuable assessments for generative models, yet gathering human feedback at scale is prohibitively costly. A practical approach to reduce the time and cost is to leverage AI feedback from VLMs [43], which has been shown to modestly align with human judgment on video quality [44, 45, 46] (see Appendix K). In this work, we assume that the VLM evaluation works as an oracle, and we align model outputs with the preferences of VLMs, which is reasonable due to their capability and the cost to be saved. Our qualitative and quantitative evaluations also confirm that the highly rated video by VLMs is generally good for us.
However, because alignment via inference-time search requires massive reward evaluation queries, we still need to build more tractable proxy rewards that do not rely on humans or external VLM APIs. The question here is what metrics for perceptual video quality can improve the feedback from VLMs. Because the criteria of videos preferred by humans are multi-objective, maximizing a single metric may lead to undesirable generation due to over-optimization. For instance, focusing exclusively on temporal consistency or frame-by-frame quality metrics can unintentionally reduce the video’s motion magnitude (see Appendix H). In this section, we first review the possible video quality metrics (Section 4.1), evaluate the Pearson correlation between these and the VLM feedback score, and then propose a reward calibration (Section 4.2), aiming to align the existing video rewards to VLMs by considering their weighted linear combination through the brute-force search of coefficients.
Figure 3: 2D-histogram and correlation between reward functions for perceptual video quality [12] and AI feedback from Gemini [14]. A single reward (e.g., subject consistency; blue) is often not aligned well with a preference from Gemini, which happens for all the prompt sets with different dynamics grades (see Figure 13). The calibrated reward, a linear combination of perceptual metrics via brute-force search (green), achieves the best Pearson correlation coefficient in all settings (statistically significant with $p<0.01$).
@@ 4.1 Metric Reward for Perceptual Video Quality
Figure 4: The coefficient of calibrated reward $w_{i}$ with feedback from Gemini. Each set of prompts, which has a different dynamics grade, requires a distinct mixture of perceptual video qualities.
@@
Experimental Setup  We leverage Gemini-1.5 [14] and GPT-4o [13] as automated raters for generated videos. We provide a prompt and generated video as inputs, instructing VLMs to assign discrete scores (from 1 to 10) based on overall visual quality (e.g., clarity, resolution, brightness, and aesthetic appeal), the appropriateness of motion for either static or dynamic scenes, the smoothness and consistency of shapes and motions, and the degree of alignment with a prompt (see Appendix I).
We select four prompt sets from two distinct datasets (see Appendix G). DEVIL [53] classifies its prompts into five categories depending on the dynamics grade, each further divided by subject type (e.g., cat, horse, truck, nature, etc.). We focus on three of the five dynamics grades (high, medium, and static) and select one prompt randomly from each subject-subdivision within a chosen category. We also draw 30 random captions from the test split of MSRVTT [54], widely used as a video benchmark.
@@ 4.2 Correlation and Reward Calibration
Figure 3 illustrates the 2D-histogram and the corresponding correlation between each metric and feedback from Gemini. Relying on a single metric often yields low correlation, which supports the multifaceted nature of perceptual video quality. See Appendix J.1 for further results, where we can see that the relative importance of each metric depends on the dynamics grade of the prompts; in highly dynamic DEVIL-high, the dynamic degree correlates more strongly with VLMs than consistency metrics. Conversely, subject consistency and motion smoothness play more prominent roles in less-dynamic DEVIL-medium or DEVIL-static. Because the aesthetic score focuses on frame-by-frame visual quality, it tends to correlate strongly with VLM in low-motion scenarios. In high-motion scenarios, in contrast, rapid movements and frequent transitions often introduce motion blur or abrupt changes in composition, reducing the frame-level aesthetic quality and thus weakening its correlation with VLMs.
Reward calibration, a weighted linear combination of these metrics, yields the highest correlation with Gemini (Figure 3, green). We select the best coefficients among brute-force candidates, based on the correlation with Gemini, for each set of prompts with a different dynamics grade (Figure 4). Prompts with a high dynamics grade, i.e., DEVIL-high, place greater weight on the dynamic degree. In contrast, prompts that describe slight motion, i.e., DEVIL-medium and DEVIL-static, place a smaller weight on it. In addition, Appendix L presents results from best-of-64 sampling with a single metric or calibrated reward, where a single metric often leads to over-optimization. This highlights the importance of reward calibration, appropriately weighting multiple criteria, as aligning with a request in the prompt.